105.0 m apart in 4.7 s. Its speed as it passes the second point is 35.5, what are its speed at the first point?

1
Expert's answer
2022-04-04T14:42:51-0400

"v=\frac s{t_2}\cdot \frac{t_1}{t_2-t_1}=24~\frac ms."